# An interpretable family of projected normal distributions and a related copula model for Bayesian analysis of hypertoroidal data

### v1

@@ Abstract
This paper introduces two families of probability distributions for Bayesian analysis of hypertoroidal data. The first family consists of symmetric distributions derived from the projection of multivariate normal distributions under specific parameter constraints. This family is closed under marginalization and hence any marginal distribution belongs to a lower-dimensional case of the same family. In particular the univariate marginal of the family is the unimodal case of the projected normal distribution on the circle. The second family is a flexible extension of the copula case of the first family, which can accommodate any univariate marginal distributions. Unlike existing models derived via projection, both families have the common advantage that their parameters possess a clear and intuitive interpretation. In addition, Markov Chain Monte Carlo algorithms are presented for Bayesian estimation of both families and a simulation study is used to demonstrate their performance. As a real data example, a meteorological data set is analyzed.
Keywords: Circular data; Circular dependence; Interpretable model.
@@ 1 Introduction
Data recorded as angles frequently appear in various academic fields. Such examples include wave directions in environmental science, directions of animal movement in biology, dihedral angles in protein backbone structures in bioinformatics, and times of crimes in political science. In practice, angular observations are often jointly observed with other angular observations, forming what is referred to as hypertoroidal data. For example, conformational angles in protein structure consist of two dihedral angles and one torsion angle of the side chain, and wave directions are typically observed at multiple locations in the sea. It is well-known that ordinary statistical methods for linear data cannot be directly applied to such hypertoroidal data and new statistical methods are needed for their analysis. See, for example, Mardia and Jupp, (1999) and Ley and Verdebout, (2017) for the general review of statistics of hypertoroidal data.
@@
A major problem in analyzing general-dimensional toroidal data is the lack of tractable and interpretable probability distributions that can serve as a foundation for parametric statistical analysis. However, if the analysis is restricted to 2- or 3-dimensional toroidal cases, there are a few distributions that are both tractable and easily interpreted. The well-known families of distributions for the 2-dimensional case are the bivariate von Mises distributions (Mardia,, 1975; Singh et al.,, 2002; Mardia et al.,, 2007; Kent et al.,, 2008; Mardia,, 2010) and copula-based models (Wehrly and Johnson,, 1980; Shieh and Johnson,, 2005; Shieh et al.,, 2011; Jones et al.,, 2015; Kato and Pewsey,, 2015). A 3-dimensional extension of Wehrly and Johnson, (1980) using the wrapped Cauchy-type copula density was recently proposed by Kato et al., (2024). Several distributions on the general-dimensional torus have been proposed in the literature. Mardia et al., (2008) and Mardia and Voss, (2014) considered a multivariate extension of the von Mises distribution, and Navarro et al., (2017) discussed its generalization. Using transformations of the multivariate normal distribution, the multivariate wrapped distribution (Jona-Lasinio et al.,, 2012), the projected normal (PN) distribution (Mastrantonio,, 2018), and the inverse stereographic normal distribution (Selvitella,, 2019) have been proposed. A tree-structured model on the general-dimensional torus was proposed by Leguey et al., (2016) in machine learning. Models on the hypertorus using other approaches include the models of Fernández-Durán and Gregorio-Domínguez, (2014) and Kim et al., (2016). Note also that, apart from the models of Leguey et al., (2016), Navarro et al., (2017) and Mastrantonio, (2018), Bayesian analysis of these multivariate distributions has not been sufficiently explored. However, these existing distributions for general-dimensional toroidal data have some limitations, such as unclear parameter interpretations, multimodality, estimations, or limited dependence structure.
@@
In this paper we propose two families of probability distributions for the Bayesian analysis of general-dimensional toroidal data. The first family, the toroidal projected normal (TPN) distribution, is derived by projecting multivariate normal distributions under specific parameter constraints. The TPN model inherits several tractable properties of the multivariate normal distribution. For example, the TPN distribution is symmetric on the torus and, in addition, the family of TPN distributions is closed under marginalization. In particular, the univariate marginal of the TPN distribution is the symmetric and unimodal case of the projected normal (PN) distribution (see, e.g., Mardia and Jupp,, 1999, Equation (3.5.49)). The second family proposed in the paper is an extension of the copula case of the TPN (CTPN) family. The CTPN is a flexible family of distributions with arbitrary univariate marginal distributions and it can be thought as a multivariate extension of the bivariate model of Wehrly and Johnson, (1980). The modes of the joint CTPN density can be expressed in a simple form under certain parameter conditions. Even though a closed-form expression for the copula density is not available in the general case, we derive it in closed form for the bivariate case. Given that the TPN, and hence also its copula, is closed under marginalization, this shows that regardless of the data dimension, the two-dimensional copula, which we know in closed form, represents the exact bivariate dependence we are assuming. For the TPN family, the parameters have clear and intuitive interpretations: each marginal mean direction and concentration is ruled by a specific parameter, and we have a parameter to model each correlation between pairs of variables. For the CTPN families, the mean directions and concentrations of the univariate marginals are determined by the parameters of the marginal distributions, while the dependence among circular variables is governed by the copula density. Similar to the TPN family, the copula parameters that rule the dependence, have clear and intuitive interpretations. Furthermore, due to their connection with multivariate normal distributions, both proposed families are well-suited for Bayesian estimation. We define a new prior distribution for the covariance matrix of the underlying normal distribution, needed for the constrains that must be imposed to obtain interpretable models, and we develop Markov chain Monte Carlo (MCMC) algorithms for Bayesian estimation of the proposed families. With a simulation study, we show that the MCMC algorithm is able to retrieve the values of the parameters used to simulate the data for both the TPN and a copula case. The proposed distributions are estimated on real meteorological data, where time series of four wind direction and two wave direction measurements are recorded.
It should be noted that the TPN is a submodel of the circular-linear distribution of Mastrantonio, (2018). The hypertoroidal submodel of Mastrantonio, (2018) offers greater flexibility than our TPN distribution, as it allows for multimodality and non-symmetry. However, as with other projected distributions (see Watson,, 1983, Section 3.6; Wang and Gelfand,, 2014; Mastrantonio et al.,, 2016), this flexibility comes at the cost of parameter interpretability. On the other hand, a tractable, interpretable, and flexible distribution can serve as a solid foundation for parametric statistical analysis. Moreover, only by constraining the parameters of the PN, it is possible to obtain a copula model with a closed-form marginal bivariate density.
The paper is organized as follows: in Section 2, we define the family of TPN distributions and provide its basic properties; in Section 3, we propose the copula version of the TPN distribution and its extension with arbitrary univariate marginal distributions; Section 4 discusse the circular correlation of our proposals; Section 5 presents the MCMC algorithm for the TPN distribution and its related copula model; in Section 6, we analyze a real dataset using our proposals; Section 7 concludes the paper. The proofs of the theoretical claims and the simulation experiments are provided in the Online Material, available on the journal website. The code used to produce the results of the paper is available at https://github.com/GianlucaMastrantonio/toroidal_projected_normal.
@@ 2.1 Idea behind the derivation of the proposed family
Before introducing the proposed family and its properties, we first explain the idea behind its derivation using the univariate and bivariate settings. In the univariate setting, a circular variable $\Theta$ is said to follow a projected normal (PN) distribution, denoted by $\mathrm{PN}((\eta_{c},\eta_{s})^{\top},\bm{\Psi})$, if a bivariate random vector $(X_{c},X_{s})^{\top}$ follows a bivariate normal distribution $N_{2}((\eta_{c},\eta_{s})^{\top},\bm{\Psi})$ and $\Theta$ is defined through the transformation:
$X_{c}=R\cos\Theta,\quad X_{s}=R\sin\Theta.$ (1)
The angle, which is the circular variable of interest, can be expressed as
@@
where $\text{atan}^{*}(y/x)$ is the modified inverse tangent function defined as an angle in $\pi,\pi)$ whose sine and cosine are given by $y/\{x^{2}+y^{2}\}^{1/2}$ and $x/\{x^{2}+y^{2}\}^{1/2}$, respectively. It is known that when $\bm{\Psi}=\mathbf{I}_{2}$ (Mardia and Jupp,, 1999, Equation (3.5.49)), the PN is symmetric around the mean direction, which is given by
$\mu=\text{atan}^{*}\left(\frac{\eta_{s}}{\eta_{c}}\right),$ (3)
and the circular concentration
$\kappa=\sqrt{\eta_{c}^{2}+\eta_{s}^{2}}$ (4)
@@
is a function of the mean vector $(\eta_{c},\eta_{s})^{\top}$. The starting point of our proposal, which to the best of our knowledge has never been pointed out before, is that, under the assumption $\bm{\Psi}=\mathbf{I}_{2}$, we can construct a PN-distributed variable by directly using $(\mu,\kappa)$ as parameters. This is possible if we assume that $(X_{c},X_{s})^{\top}\sim N((\kappa,0)^{\top},\mathbf{I}_{2})$, with
$X_{c}=R\cos(\Theta-\mu),\quad X_{s}=R\sin(\Theta-\mu).$ (5)
@@
We can now move to the bivariate setting. If we have two circular variables $\Theta_{1}$ and $\Theta_{2}$, which are derived from the linear variables
$\bm{X}_{c}=\begin{pmatrix}X_{c,1}\\ X_{c,2}\end{pmatrix}=\begin{pmatrix}R_{1}\cos(\Theta_{1}-\mu_{1})\\ R_{2}\cos(\Theta_{2}-\mu_{2})\end{pmatrix}\quad\mbox{and}\quad\bm{X}_{s}=\begin{pmatrix}X_{s,1}\\ X_{s,2}\end{pmatrix}=\begin{pmatrix}R_{1}\sin(\Theta_{1}-\mu_{1})\\ R_{2}\sin(\Theta_{2}-\mu_{2})\end{pmatrix}$, (6)
@@
an easy way to introduce dependence between the circular variables while keeping the marginal distribution invariant (see Theorem 2 below for a formal proof) is to let $\bm{X}_{c}\perp\bm{X}_{s}$ and assume
$\bm{X}_{c}\sim N_{d}(\bm{\kappa},\bm{\Sigma}_{c}),\quad\bm{X}_{s}\sim N_{d}(\bm{0},\bm{\Sigma}_{s}),$ (7)
@@
where $\bm{\kappa}\in(\mathbb{R}^{+})^{2}$, and both $\bm{\Sigma}_{c}$ and $\bm{\Sigma}_{s}$ are correlation matrices. Depending on the values of $\rho_{c}=[\bm{\Sigma}_{c}]_{1,2}$ and $\rho_{s}=[\bm{\Sigma}_{s}]_{1,2}$, we may observe different types of dependence. For interpretability purposes and parsimony of parameters, we want to have a single parameter $\rho$ such that, for positive values, the two circular variables tend to rotate in the same direction, while for negative values, they tend to rotate in opposite directions.
@@
From Wang and Gelfand, (2013) we know that $E(\sin(\Theta_{1}-\mu_{1}))=E(\sin(\Theta_{2}-\mu_{2}))=0$. It is evident that $X_{s,j}$ and $\sin(\Theta_{1}-\mu_{1})$ have the same sign and, moreover, the larger the value of $X_{s,j}$, the larger $\sin(\Theta_{1}-\mu_{1})$ is. Hence, a positive $\rho_{s}$ will induce a positive correlation between the two sine transformations while a negative one will have the opposite effect. To induce the correct correlation between the two circular variables, we also need to define the correlation between the two $X_{c,j}$, which induces a dependence between the cosines. In this case, regardless of whether the relationship between $\sin(\Theta_{1}-\mu_{1})$ and $\sin(\Theta_{2}-\mu_{2})$ is proportional or inversely proportional, as $\Theta_{j}-\mu_{j}\in[-\pi,\pi)$ decreases within $\pi,0]$ or increases within $0,\pi)$, $\cos(\Theta_{j}-\mu_{j})$ decreases, as does $X_{c,j}$. This suggests that, for the sign of $\rho_{c}$ to reflect the relationship between $\Theta_{1}$ and $\Theta_{2}$, the parameter $\rho_{c}$ must always be positive. For this reason, we assume $\bm{\Sigma}_{s}]_{1,2}=\rho$ and $\bm{\Sigma}_{c}]_{1,2}=|\rho|$, which is the basic idea of our proposal. A generalization to arbitrary dimension is proposed below.
@@ 2.2 Definition and basic properties
Let $\bm{X}_{c}$ and $\bm{X}_{s}$ be two independent $d$-dimensional random vectors distributed as $N_{d}(\bm{\kappa},\bm{\Sigma}_{c})$ and $N_{d}(\bm{0},\bm{\Sigma})$, respectively, where $\bm{\kappa}=(\kappa_{1},\ldots,\kappa_{d})^{\top}$, $\kappa_{1},\ldots,\kappa_{d}\geq 0$, and $\bm{\Sigma}_{c}$ and $\bm{\Sigma}$ are correlation matrices with elements $\bm{\Sigma}]_{j,k}=\rho_{jk}$ and $\bm{\Sigma}_{c}]_{j,k}=|\rho_{jk}|$. Transform $\bm{X}_{c}$ and $\bm{X}_{s}$ into
$\bm{X}_{c}=\begin{pmatrix}R_{1}\cos(\Theta_{1}-\mu_{1})\\ R_{2}\cos(\Theta_{2}-\mu_{2})\\ \vdots\\ R_{d}\cos(\Theta_{d}-\mu_{d})\end{pmatrix}\quad\mbox{and}\quad\bm{X}_{s}=\begin{pmatrix}R_{1}\sin(\Theta_{1}-\mu_{1})\\ R_{2}\sin(\Theta_{2}-\mu_{2})\\ \vdots\\ R_{d}\sin(\Theta_{d}-\mu_{d})\end{pmatrix}$, (8)
with $R_{j}>0$, $\Theta_{j}\in[-\pi,\pi)$, and $\pi\leq\mu_{1},\ldots,\mu_{d}<\pi$. Then the random vector $(\Theta_{1},\ldots,\Theta_{d})$ consists of $d$-dimensional circular variables living on the hypertorus $\pi,\pi)^{d}$, and we refer to their joint distribution as the (hyper-)toroidal projected normal (TPN) distribution.
For convenience, we write $(\Theta_{1},\ldots,\Theta_{d})\sim TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$, with $\bm{\mu}=(\mu_{1},\ldots,\mu_{d})^{\top}$, if the random vector $(\Theta_{1},\ldots,\Theta_{d})$ follows the distribution in Definition 1. It should be noted that both $\bm{\Sigma}$ and $\bm{\Sigma}_{c}$ are positive definite correlation matrices. However, not every matrix $\bm{\Sigma}$ produces a matrix $\bm{\Sigma}_{c}$ that is positive definite. This poses challenges for the estimation of the covariance matrices since we need to restrict the space where $\bm{\Sigma}$ is defined. We will tackle this issue with an appropriate prior distribution in Section 5, where the MCMC algorithm is presented.
From the definition is easy to prove the following Theorem:
@@
Let $\bm{\Theta}\sim TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$. Then the distribution of $\bm{\Theta}$ is symmetric about $\bm{\mu}$, namely, $\bm{\Theta}-\bm{\mu}\stackrel{{\scriptstyle d}}{{=}}\bm{\mu}-\bm{\Theta}$.
The symmetry of our model leads to clear interpretation of the parameters, which include not only $\bm{\mu}$ itself but also $\bm{\kappa}$ and $\bm{\Sigma}$; see Theorem 3 and Figures 1 and 2 below.
@@
Let $(\Theta_{1},\ldots,\Theta_{d})\sim TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$. Suppose $\bm{\Theta}_{1}=(\Theta_{1},\ldots,\Theta_{d_{1}})$ and $\bm{\Theta}_{2}=(\Theta_{d_{1}+1},\ldots,\Theta_{d})$. Then
$\bm{\Theta}_{1}\sim TPN_{d_{1}}(\bm{\mu}_{1},\bm{\kappa}_{1},\bm{\Sigma}_{11})\quad\text{and}\quad\bm{\Theta}_{2}\sim TPN_{d_{2}}(\bm{\mu}_{2},\bm{\kappa}_{2},\bm{\Sigma}_{22}),$
@@
Note that any permutation of the random vector $(\Theta_{1},\ldots,\Theta_{d})$ also has the TPN distribution due to the symmetric form of the distribution and is hence closed under marginalization.
@@
Let $(\Theta_{1},\ldots,\Theta_{d})\sim TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$. Then the marginal distribution of $\Theta_{j}$ $(j=1,\ldots,d)$ is a symmetric and unimodal case of the PN distribution, with density
$f(\theta_{j})=\frac{1}{\sqrt{2\pi}}\phi(\kappa_{j})+\kappa_{j}\cos(\theta_{j}-\mu_{j})\phi(\kappa_{j}\sin(\theta_{j}-\mu_{j}))\Phi(\kappa_{j}\cos(\theta_{j}-\mu_{j})),$ (9)
@@
The proof for Theorem 3 is straightforward and follows directly from Mardia and Jupp, (1999, Section 3.5.6): it is therefore omitted.
From Theorems 13, it follows that our model can be considered a toroidal extension of the symmetric and unimodal PN distribution on the circle. Theorem 2 also gives the distribution of any possible combination of two circular variables, which are marginally TPN distributed. Then, given the discussion in Section 2.1, we can clearly see that $\mu_{j}$ is the mean direction of $\Theta_{j}$, the concentration of $\Theta_{j}$ depends only on $\kappa_{j}$, while the magnitude and sign of $\rho_{jk}$ measure the dependence between $\Theta_{j}$ and $\Theta_{k}$. While the TPN distribution is a special case of the poly-cylindrical distribution of Mastrantonio, (2018), our model, in comparison with Mastrantonio, (2018), has the advantage that each parameter can be clearly interpreted. Plots of the univariate and bivariate densities of the TPN are shown in Figures 1 and 2. The densities in Figure 1 are computed using (9), while those in Figure 2 are computed via numerical integration of the expression (12) given later.

### v2

@@ Abstract
This paper introduces two families of probability distributions for Bayesian analysis of hypertoroidal data. The first family consists of symmetric distributions derived from the projection of multivariate normal distributions under specific parameter constraints. This family is closed under marginalization and hence any marginal distribution belongs to a lower-dimensional case of the same family. In particular the univariate marginal of the family is the unimodal case of the projected normal distribution on the circle. The second family is a flexible extension of the copula case of the first family, which can accommodate any univariate marginal distributions. Unlike existing models derived via projection, both families have the common advantage that their parameters possess a clear and intuitive interpretation. The use of latent variables simplifies Bayesian estimation using Markov chain Monte Carlo algorithms. The usefulness of the proposed families is demonstrated through the analysis of a meteorological dataset.
Keywords: Circular data; Circular dependence; Directional statistics; Interpretable parameter; Unimodality.
@@ 1 Introduction
Data recorded as angles frequently appear in various academic fields. Such examples include wave directions in environmental science, directions of animal movement in biology, dihedral angles in protein backbone structures in bioinformatics, and times of crimes in political science. In practice, angular observations are often jointly observed with other angular observations, forming what is referred to as hypertoroidal data. For example, conformational angles in protein structure consist of two dihedral angles and one torsion angle of the side chain, and wave directions are typically observed at multiple locations in the sea. It is well-known that ordinary statistical methods for linear data cannot be directly applied to such hypertoroidal data and new statistical methods are needed for their analysis. See, for example, Mardia1999a, Ley2017a and Pewsey2021 for the general review of statistics of hypertoroidal data.
@@
A major problem in analyzing hypertoroidal data is the lack of tractable and interpretable probability distributions that can serve as a foundation for parametric statistical analysis. This contrasts with the analysis of hyperspherical data, for which a variety of such distributions are available; see, e.g., Scealy2019, Kato2020 and Pewsey2021 for recent reviews. However, if the analysis is restricted to 2- or 3-dimensional toroidal cases, there are a few distributions that are both tractable and easily interpreted. The well-known families of distributions for the 2-dimensional case include the bivariate von Mises distributions (Mardia1975b; Singh2002; Mardia2007; Kent2008; Mardia2010) and copula-based models (Wehrly1980; Shieh2005; Shieh2011; Jones2015; Kato2015a). A 3-dimensional extension of Wehrly1980 using the wrapped Cauchy-type copula density was recently proposed by Kato2024. Several distributions on the general-dimensional torus have been proposed in the literature. Mardia2008 and Mardia2014 considered a multivariate extension of the von Mises distribution, and Navarro2017 discussed its generalization. The inverse stereographic normal distribution has been proposed by Selvitella2019. A tree-structured model on the general-dimensional torus was proposed by Leguey2016 in machine learning. Models on the hypertorus using other approaches include the models of Fernandez-Duran2014 and Kim2016a. However, existing distributions for general-dimensional toroidal data have certain limitations, such as unclear parameter interpretations, multimodality, and/or a limited range of attainable dependence strength between variables. In addition, a common issue is that parameter estimation via the maximum likelihood approach is often difficult due to a complicated form of the probability density function.
Recently, Bayesian approaches have been increasingly adopted for modeling directional data, particularly in relation to spatial and spatio-temporal directional data analysis. Many of these successful approaches are based on projections or wrappings of multivariate normal distributions; see, e.g., Jona-Lasinio2012, Wang2013a; Wang2014, Wang2015, Mastrantonio2016a, and Hernandez-Stumpfhauser2017. Although these models generally do not have closed-form expressions for their probability density functions, they provide elegant solutions for Bayesian estimation by introducing latent variables. However, these models are not designed for hypertoroidal data, which are particularly challenging to handle using frequentist approaches, except for a special case of the model proposed by Mastrantonio2018, which nonetheless suffers from issues such as multimodality, unclear parameter interpretations, and a limited range of dependence between variables.
@@
In this paper we propose two families of probability distributions for the Bayesian analysis of hypertoroidal data. The first family, the (hyper-)toroidal projected normal (TPN) distribution, is derived by projecting multivariate normal distributions under specific parameter constraints. The TPN model inherits various tractable properties of the multivariate normal distribution. For example, unlike most of existing projected distributions for directional data, the TPN distribution is symmetric around the mean direction. The family of TPN distributions is closed under marginalization and, in particular, the univariate marginal of the TPN distribution corresponds to the symmetric and unimodal case of the projected normal (PN) distribution (see, e.g., Mardia1999a, Equation (3.5.49)). The second family proposed in this paper is an extension of the copula case of the TPN (CTPN) family. The CTPN is a flexible family of distributions with arbitrary univariate marginal distributions. The bivariate case of the CTPN belongs to the well-known family of Wehrly1980 and therefore has a simple expression for its density. The modes of the joint CTPN density can be expressed in a simple form under certain parameter conditions. The CTPN can describe a wide range of dependence structures between variables, including independence as well as perfect positive and negative dependence. For both the TPN and CTPN families, the parameters have clear and intuitive interpretations. For the TPN family, each marginal mean direction and concentration is governed by a specific parameter, and an additional parameter controls the correlation between each pair of variables. For the CTPN family, each marginal distribution is specified freely, with its mean direction and other characteristics determined by the parameters of the marginal density, while the dependence among circular variables is governed by the copula density. As with existing projected distributions, the proposed families generally do not have closed-form expressions for their densities. However, the introduction of latent variables greatly simplifies Bayesian estimation for both families. We define a new prior distribution for the covariance matrix of the underlying normal distribution, needed for the constrains that must be imposed to obtain interpretable models, and we develop Markov chain Monte Carlo (MCMC) algorithms for Bayesian estimation of the proposed families. Through a simulation study, it is seen that the MCMC algorithm successfully recover the values of the parameters used to simulate the data for both the TPN and the copula case. We apply the proposed distributions to real meteorological data consisting of time series of four wind-direction and two wave-direction measurements.
The code used to produce the results of the real data example in §6 and the simulation examples in the Supplementary Material is available at https://github.com/GianlucaMastrantonio/toroidal_projected_normal.
@@ 2.1 Idea behind the derivation of the proposed family
Before introducing the proposed family and its properties, we first explain the idea behind its derivation using the univariate and bivariate settings. In the univariate setting, a circular variable $\Theta$ is said to follow a projected normal (PN) distribution, denoted by $\mathrm{PN}((\eta_{c},\eta_{s})^{\top},\bm{\Psi})$, if a bivariate random vector $(X_{c},X_{s})^{\top}$ follows a bivariate normal distribution ${\cal N}_{2}((\eta_{c},\eta_{s})^{\top},\bm{\Psi})$, and $\Theta$ is defined through the transformation:
@@
where $\text{atan}^{*}(y/x)$ denotes the modified arctangent function that returns an angle in $\pi,\pi)$ for $(x,y)\neq(0,0)$, whose sine and cosine are $y/\{x^{2}+y^{2}\}^{1/2}$ and $x/\{x^{2}+y^{2}\}^{1/2}$, respectively (see Mardia1999a, Equation (2.2.4)). It is known that when $\bm{\Psi}=\bm{I}_{2}$, the PN is symmetric around the mean direction $\mu=\text{atan}^{*}(\eta_{s}/\eta_{c}),$ and the concentration parameter $\kappa=\{\eta_{c}^{2}+\eta_{s}^{2}\}^{1/2}$ is a function of the mean vector $(\eta_{c},\eta_{s})^{\top}$ (Mardia1999a, Equation (3.5.49)). The starting point of our proposal, which to the best of our knowledge has never been pointed out before, is that, under the assumption $\bm{\Psi}=\bm{I}_{2}$, we can construct a PN-distributed variable by directly using $(\mu,\kappa)$ as parameters. This is possible if we assume that $(X_{c},X_{s})^{\top}\sim{\cal N}_{2}((\kappa,0)^{\top},\bm{I}_{2})$, with $\Theta=\mu+\text{atan}^{*}(X_{s}/X_{c}).$
@@
We can now move to the bivariate setting. If we derive two circular variables $\Theta_{1}$ and $\Theta_{2}$ from the linear variables $\{X_{c,1},X_{c,2},X_{s,1},X_{s,2}\}$ via
$\Theta_{1}=\mu_{1}+\text{atan}^{*}\left(\frac{X_{s,1}}{X_{c,1}}\right),\quad\Theta_{2}=\mu_{2}+\text{atan}^{*}\left(\frac{X_{s,2}}{X_{c,2}}\right),$ (2)
@@
a simple way to introduce dependence between $\Theta_{1}$ and $\Theta_{2}$ while keeping the marginal distribution invariant (see Theorem 2 below for a formal proof) is to let $(X_{c,1},X_{c,2})$ and $(X_{s,1},X_{s,2})$ be independent and assume
$\begin{pmatrix}X_{c,1}\\ X_{c,2}\end{pmatrix}\sim{\cal N}_{2}(\bm{\kappa},\bm{\Sigma}_{c}),\quad\begin{pmatrix}X_{s,1}\\ X_{s,2}\end{pmatrix}\sim{\cal N}_{2}(\bm{0},\bm{\Sigma}_{s}),$ (3)
@@
where $\bm{\kappa}\in(\mathbb{R}^{+})^{2}$, and both $\bm{\Sigma}_{c}$ and $\bm{\Sigma}_{s}$ are correlation matrices. Depending on the values of $\rho_{c}=[\bm{\Sigma}_{c}]_{1,2}$ and $\rho_{s}=[\bm{\Sigma}_{s}]_{1,2}$, different types of dependence may be observed, where $\bm{\Sigma}]_{j,k}$ denotes the $(j,k)$ element of $\bm{\Sigma}$. For interpretability and parameter parsimony, we aim to introduce a single parameter $\rho$ such that, for positive values, the two circular variables tend to be positively related, whereas for negative values, they tend to be negatively related.
@@
From Wang2013a, we know that $E\{\sin(\Theta_{1}-\mu_{1})\}=E\{\sin(\Theta_{2}-\mu_{2})\}=0$. It is evident that $X_{s,j}$ and $\sin(\Theta_{j}-\mu_{j})$ have the same sign and, moreover, the larger the value of $X_{s,j}$, the larger the value of $\sin(\Theta_{j}-\mu_{j})$. Hence, a positive $\rho_{s}$ induces a positive correlation between the two sine transformations and hence between $\Theta_{1}-\mu_{1}$ and $\Theta_{2}-\mu_{2}$, whereas a negative $\rho_{s}$ has the opposite effect. To induce the appropriate correlation between the two circular variables, we also need to define the correlation between the two $X_{c,j}$’s which induces a dependence between the cosines. In this case, regardless of whether the relationship between $\Theta_{1}-\mu_{1}$ and $\Theta_{2}-\mu_{2}$ is positive or negative, it is natural to assume that $\cos(\Theta_{j}-\mu_{j})$ increases with $X_{c,j}$ for $j=1,2$, due to the identity $\cos(\Theta_{j}-\mu_{j})=\cos\{-(\Theta_{j}-\mu_{j})\}$. This suggests that, for the sign of $\rho_{c}$ to reflect the relationship between $\Theta_{1}$ and $\Theta_{2}$, the parameter $\rho_{c}$ should always be positive. For this reason, we set $\bm{\Sigma}_{s}]_{1,2}=\rho$ and $\bm{\Sigma}_{c}]_{1,2}=|\rho|$, which is the basic idea of our proposal. A generalization to arbitrary dimensions is proposed below.
@@ 2.2 Definition and basic properties
Let $\bm{X}_{c}=(X_{c,1},\ldots,X_{c,d})^{\top}$ and $\bm{X}_{s}=(X_{s,1},\ldots,X_{s,d})^{\top}$ be two independent $d$-dimensional random vectors distributed as ${\cal N}_{d}(\bm{\kappa},\bm{\Sigma}_{c})$ and ${\cal N}_{d}(\bm{0},\bm{\Sigma})$, respectively, where $\bm{\kappa}=(\kappa_{1},\ldots,\kappa_{d})^{\top}$, $\kappa_{1},\ldots,\kappa_{d}\geq 0$, and $\bm{\Sigma}_{c}$ and $\bm{\Sigma}$ are correlation matrices with the $(j,k)$ elements $\bm{\Sigma}]_{j,k}=\rho_{jk}$ and $\bm{\Sigma}_{c}]_{j,k}=|\rho_{jk}|$ $(1\leq j,k\leq d)$. Define the circular random variables
$\Theta_{1}=\mu_{1}+\text{atan}^{*}\left(\frac{X_{s,1}}{X_{c,1}}\right),\quad\ldots,\quad\Theta_{d}=\mu_{d}+\text{atan}^{*}\left(\frac{X_{s,d}}{X_{c,d}}\right),$ (4)
with $\Theta_{j}\in[-\pi,\pi)$ and $\mu_{j}\in[-\pi,\pi)$. Then the random vector $(\Theta_{1},\ldots,\Theta_{d})$ consists of $d$-dimensional circular variables taking values on the hypertorus $\pi,\pi)^{d}$, and we refer to their joint distribution as the (hyper-)toroidal projected normal (TPN) distribution.
As is clear from the definition, random variates from the TPN distributions can be efficiently generated using those from multivariate normal distributions.
For convenience, we write $(\Theta_{1},\ldots,\Theta_{d})\sim\mathrm{TPN}_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$, with $\bm{\mu}=(\mu_{1},\ldots,\mu_{d})^{\top}$, if the random vector $(\Theta_{1},\ldots,\Theta_{d})$ follows the distribution in Definition 1. For any $\bm{\Sigma}$, $\bm{\Sigma}_{c}$ is always positive definite when $d=2$, but this is not generally the case when $d\geq 3$. Nevertheless, our numerical simulations suggest that when the positive definite matrices $\bm{\Sigma}$ with $d\geq 3$ are generated from Inverse Wishart distributions, the corresponding $\bm{\Sigma}_{c}$ is often positive definite as well. In Bayesian estimation, appropriate prior distributions must be specified for both $\bm{\Sigma}$ and $\bm{\Sigma}_{c}$, see §5.
Basic properties of the TPN family are provided below. Here and throughout this section, assume that $\bm{\Theta}=(\Theta_{1},\ldots,\Theta_{d})^{\top}\sim\mathrm{TPN}_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$.
@@
The distribution of $\bm{\Theta}$ is symmetric about $\bm{\mu}$, namely, $\bm{\Theta}-\bm{\mu}\stackrel{{\scriptstyle d}}{{=}}\bm{\mu}-\bm{\Theta}$.
The symmetry of our model leads to clear interpretation of the parameters, which include not only $\bm{\mu}$ itself but also $\bm{\kappa}$ and $\bm{\Sigma}$; see Theorem 3 and Figs. 1 and 2 below.
@@
Let $\bm{\Theta}_{1}=(\Theta_{1},\ldots,\Theta_{d_{1}})$ and $\bm{\Theta}_{2}=(\Theta_{d_{1}+1},\ldots,\Theta_{d})$. Then
$\bm{\Theta}_{1}\sim\mathrm{TPN}_{d_{1}}(\bm{\mu}_{1},\bm{\kappa}_{1},\bm{\Sigma}_{11})\quad\text{and}\quad\bm{\Theta}_{2}\sim\mathrm{TPN}_{d_{2}}(\bm{\mu}_{2},\bm{\kappa}_{2},\bm{\Sigma}_{22}),$
@@
Any permutation of the random vector $(\Theta_{1},\ldots,\Theta_{d})$ also has the TPN distribution due to the symmetric form of the distribution and is hence closed under marginalization.
@@
The marginal distribution of $\Theta_{j}$ $(j=1,\ldots,d)$ is a symmetric and unimodal case of the PN distribution with density
$f(\theta_{j})=\frac{1}{\surd(2\pi)}\phi(\kappa_{j})+\kappa_{j}\cos(\theta_{j}-\mu_{j})\phi\{\kappa_{j}\sin(\theta_{j}-\mu_{j})\}\Phi\{\kappa_{j}\cos(\theta_{j}-\mu_{j})\},$ (5)
@@
The proof for Theorem 3 is straightforward and follows directly from Mardia1999a: it is therefore omitted.
From Theorems 13, it follows that our model can be considered a toroidal extension of the symmetric and unimodal PN distribution on the circle. Theorem 2 also gives the distribution of any possible combination of two circular variables, which are marginally TPN distributed. Then, given the discussion in §2.1, we can clearly see that $\mu_{j}$ is the mean direction of $\Theta_{j}$, the concentration of $\Theta_{j}$ depends only on $\kappa_{j}$, while the magnitude and sign of $\rho_{jk}$ measure the dependence between $\Theta_{j}$ and $\Theta_{k}$.